OpenStudy (anonymous):

What value completes the square for the expression? x2 - 18x

OpenStudy (anonymous):

@ajeetkumarseo

OpenStudy (anonymous):

@wolf1728

OpenStudy (wolf1728):

x² -18x = 0 To complete the square, you first divide the equation by the coefficient of x². There is no coefficient of x² so this step is done. Move the non-X term to the right. There is no non-X term so this step is also done. Now we must take the coefficient of x divide it by 2 square that then add that to both sides of the equation. Coefficient of x =-18 -18 / 2 = -9 -9² = 81 Now we add that to both sides of the equation: x² -18x +81= 81 Then take the square root of both sides. (x-9) = 9 x = 18 The square root of 81 can also be -9 so the other 'x' value is: (x-9) = -9 x = 0 Web page for completing the square: http://www.1728.org/quadr2.htm